# Erasing Without Remembering: Implicit Knowledge Forgetting in Large Language Models

### v2

@@ 5.2 Perturbed Distribution Matching
Building on MSM, we identify the top-$K$most sensitive tokens and inject random noise into their embeddings before feeding them into the model, resulting in the perturbed unlearning sample $x^{\prime}$. We first pass the original unlearning sample $x$ into the unlearning model $f_{\theta_{u}}$ to obtain the clean-run next-token probability distribution $p(y|x)$, where $y=f_{\theta_{u}}(y|x)$ represents the model’s output. Similarly, the corrupted-run next-token probability distribution $p(y|x^{\prime})$ is obtained by feeding the perturbed sample $x^{\prime}$ into $f_{\theta_{u}}$. As illustrated in Fig. 3, the clean-run probability distribution $p(y|x)$ assigns the highest probabilities to tokens that are informative and fact-related. In contrast, for the corrupted-run distribution $p(y|x^{\prime})$, the model fails to recall related facts, resulting in token generation that relies primarily on context or grammar, with top-ranked tokens being fact-unrelated.
@@
1:  Identify the top-$K$most sensitive tokens in $x$ using MSM.
2:  Introduce random noise to the $K$tokens embeddings to obtain the perturbed sample: $x^{\prime}$.
@@
5:  Subtract the clean-run distribution from the corrupted-run distribution to emulate forgetting:$p(Y_{t}|y_{<t})=p(y|x^{\prime})-C\cdot p(y|x).$
6:  Fine-tune $f_{\theta_{u}}$ to match $p(Y_{t}|y_{<t})$ by minimising the step-wise KL divergence:$L=-\sum_{i=1}^{t}\sum_{Y_{i}\in V}p(Y_{i}|y_{<i})\log q_{\theta}(Y_{i}|y_{<i}),$ where $V$ is the vocabulary and $q_{\theta}$ represents the predicted distributions of $f_{\theta_{u}}$.
@@
6 Evaluation on UGBench
In this section, we evaluate PerMU on UGBench and analyze the results. The implementation parameters of PerMU are consistent with those of the baselines. We integrate GDR with $RW=1$, set the percentage of perturbed tokens to $K=0.4$, and maintain perturbation ratio $P=0.4$, coefficient $C=0.1$.
@@ 6.1 Improved Unlearning Capabilities
When tested on TOFU Forget05 dataset using LLaMA2-7B, as shown in Tab. III, PerMU demonstrates remarkable effectiveness in unlearning knowledge. Specifically, PerMU outperforms the baselines with a relative improvement of up to 50.40% (78.67 $\rightarrow$39.02) in Forget Probability and 11.3% (56.27$\rightarrow$62.63) on Forget Truth Ratio.
@@
TABLE VI: Experimental results for inverted relation data, subject-replaced data, and one-hop reasoned data on the ZsRE dataset using Phi-1.3B. Most methods struggle to forget the “hard” in-scope knowledge that are logically related to the original unlearning sample. Nevertheless, PerMU achieves an improvement of up to 9.41% on FRT ratio (0.85$\rightarrow$0.93).
@@ 6.2 Improved Generalisation Capabilities
When evaluated on the ZsRE dataset, as shown in Tab. VI, although the ICL-based method demonstrates superior unlearning performance, it suffers from lower model utility, which limits its applicability in real-world scenarios. Apart from it, PerMU achieves the best FRT ratio across all three subsets, particularly excelling on inverted relation data, where it shows a relative improvement of up to 9.41% (0.85$\rightarrow$0.93) on FRT ratio. This highlights PerMU’s ability to strike a better balance between unlearning and retaining knowledge. However, the absolute generalisation performance of all methods in forgetting knowledge logically related to the original unlearning samples remains suboptimal, particularly when it comes to unlearning knowledge that requires one-hop reasoning, showing a gap of up to 2.23 times lower than the Retain model. Enhancing the absolute generalisation ability of the unlearned model is crucial in machine unlearning, leaving room for further exploration by the community.
@@ 6.3 Discussions
TABLE VIII: Comparison between the average MSM values of subject words and the other words on UGBench datasets. Subject words exhibit higher MSM values than other tokens, indicating greater sensitivity.
@@ 6.4 Fast Alternative Implementation of PerMU
To reduce the computational overhead of identifying the top-$K$most sensitive tokens for each unlearning sample, we explore a more efficient implementation of PerMU. We begin by computing the MSM to determine the types of tokens to which the model is most sensitive. Experiments are conducted on UGBench datasets using Phi-1.3B.
@@
We present the experimental results of $\textsc{PerMU}^{\dagger}$ on UGBench. As shown in Tab.IV, $\textsc{PerMU}^{\dagger}$ achieves unlearning performance comparable to existing methods, even surpasses PerMU by up to 17.64% (17.91 $\rightarrow$ 14.75) on Rephrased TOFU Forget01 Probability, and 8.19 (63.09 $\rightarrow$ 68.26) on Rephrased TOFU Forget05 TR. Although $\textsc{PerMU}^{\dagger}$ underperforms PerMU on Harry Potter dataset, it still outperforms baseline methods, achieving relative improvements of up to 15.47% (87.35 $\rightarrow$ 71.88) in Forget ROUGE, 8.69% (65.05 $\rightarrow$ 56.36) in Forget Probability and 15.36% (84.90 $\rightarrow$ 69.54) in Forget F1 when evaluated on LLaMA2-7B, all while maintaining high model utility. These results demonstrate the effectiveness of $\textsc{PerMU}^{\dagger}$ in achieving efficient forgetting with minimal computational overhead.
@@ 7.1 Evaluation on Generation Quality
The impact of machine unlearning on language models is intricate, requiring a thorough and comprehensive evaluation to fully understand its effects. To this end, we perform additional tests to evaluate the generation quality of existing methods. Building on the previous work [38], we introduce the Fluency metric to measure the fluency of the unlearned model’s output sentences. Fluency is measured by the weighted average of bi- and tri-gram entropies [51], defined as $-\sum_{k}f(k)\log_{2}f(k)$, where $f(\cdot)$ represents the $n$-gram frequency distribution. A higher Fluency score indicates more informative and diverse text generation. Experiments are conducted on LLaMA2-7B across all subsets.
As shown in Tab. IX, PerMU achieves a higher Fluency ratio on most datasets. Notably, when tested on the ZsRE dataset, PerMU outperforms others by nearly twofold on both the Forget dataset and the Rephrased Forget dataset. Moreover, PerMU achieves the highest average Fluency score across all datasets, with a relative improvement of up to 37.5% (0.40$\rightarrow$0.55). We present a case study in Tab. X, where Gradient Ascent generates low-quality outputs with repetitive content. In contrast, PerMU produces more fluent and readable sentences.
@@ 7.4 Computational Overhead
We evaluated the computational overhead of existing methods from two perspectives: training step time and memory usage. The experiments were conducted on the Phi-1.3B model using the TOFU Forget01 dataset, with a batch size of 4 for 5 epochs. We report the total training step time and the maximum memory usage. As shown in Tab. XIV, although PerMU requires more training time, most of the overhead comes from computing the MSM to identify the top-$K$sensitive tokens. Notably, this step can be performed offline prior to training, significantly reducing both time and memory consumption during actual training. Moreover, our proposed fast variant, $\textsc{PerMU}^{\dagger}$, introduces no additional cost in training step time or memory usage, confirming its computational efficiency.
@@ 7.5 Percentage of Perturbed Tokens $K$
In PerMU, we identify the top-$K$ most sensitive tokens and apply perturbations to them. The choice of $K$ directly influences the unlearning effect. We analyze the impact of varying $K$ from 0.1 to 1.0 in increments of 0.1, where each value represents $K$of the tokens to be perturbed, while keeping all other hyper-parameters fixed. As shown in Tab. XIII, increasing $K$ consistently enhances the unlearning ability, with Forget Rouge improved by up to 5.62% (47.19 $\rightarrow$ 41.57) on the TOFU dataset. However, this improvement comes at the cost of reduced model utility. This trade-off is expected, as perturbing more tokens makes it harder for the model to recover and correctly predict the unlearning samples. Although the Forget Retain Trade-off (FRT) score peaks when $K=0.8$ to $0.9$, the performance on retained knowledge drops significantly—up to 5.76% in Rouge (82.04 $\rightarrow$ 76.28) and 3.84% in Probability (88.47 $\rightarrow$ 84.63). Such degradation may negatively impact the model’s utility on preserved knowledge. To strike a balance between forgetting and retention, we set $K=0.4$ in our main experiments, which achieves effective unlearning while preserving most of the model’s capacity to retain relevant knowledge.
@@ 7.6 Perturbation Ratio $P$
TABLE XVI: The impact of the Tuning Coefficient $C$. Experiments are conducted on the TOFU Forget01 dataset using Phi-1.3B. Increasing $C$ results in an improved unlearning effect, but at the cost of decreased model utility. To balance model utility with effective unlearning, we select $C=0.1$ for our experiments.
@@ 7.7 Tuning Coefficient $C$
We investigate the impact of the Tuning Coefficient $C$, which varies from $0.0$ to $1.0$ in increments of $0.1$. The experiments are conducted on the TOFU Forget01 dataset using Phi-1.3B, with all other parameters held constant. As shown in Tab. XVI, the best model utility is achieved when $C=0.0$, where only the corrupted-run probability distribution is used. While this setting maintains high model utility, the unlearning effect is insufficient, as the top-ranked token in the clean-run probability distribution is not fully suppressed. As $C$ increases, the unlearning effect improves, reaching its peak when $C=0.4$. Then it begins to fluctuate as $C$ continues to increase. Correspondingly, model utility decreases with larger values of $C$, which is expected, as higher $C$ values subtract more information from $p(y|x^{\prime})$, potentially disrupting the distribution of irrelevant knowledge. To balance model utility with effective unlearning, we select $C=0.1$ for our experiments.
7.8 Discrete-Token Level Perturbation
As shown in Tab. XVII, $\textsc{PerMU}^{\dagger}_{dis}$ exhibits exceptional unlearning capability, outperforming $\textsc{PerMU}^{\dagger}$ in Forget ROUGE across all datasets and models. Moreover, $\textsc{PerMU}^{\dagger}_{dis}$ achieves an absolute 19.26% Forget ROUGE and 4.64% Forget Probability on Forget05 when using LLaMA2-7B, surpassing $\textsc{PerMU}^{\dagger}$ by up to 11.16% (30.42$\rightarrow$19.26) and 11.65% (16.29$\rightarrow$4.64), respectively, though with a slightly lower model utility of 0.92%. However, when tested on Phi-1.3B, $\textsc{PerMU}^{\dagger}_{dis}$ does not consistently exhibit superior unlearning performance, and the model utility drops by up to 4.44% (50.94$\rightarrow$46.50) as the number of unlearning samples increases. Despite this, the FRT ratio of $\textsc{PerMU}^{\dagger}_{dis}$ still outperforms other baselines. In summary, perturbing subject words at the discrete-token level can also prevent the model from recalling the fact and generate fact-unrelated probability distributions, thus achieving unlearning. Both embedding-layer and discrete-token-level noise methods can achieve effective unlearning but result in different trade-offs. Given that the knowledge retention ability of $\textsc{PerMU}^{\dagger}_{dis}$ may decline as the amount of forgotten data increases, we choose to add noise at the embedding layer as a more promising alternative.
TABLE XVII: Experimental results for implementing perturbation at the discrete-token level to subject words, denoted as $\textsc{PerMU}_{dis}$. Surprisingly, $\textsc{PerMU}^{\dagger}_{dis}$ demonstrates a stronger unlearning effect on some datasets, particularly achieving 19.26% in Forget Rouge and 4.64% in Probability on the Forget01 dataset using LLaMA2-7B. These results highlight the adaptability of our approach across different dimensions.
@@ 7.8 Discrete-Token Level Perturbation
TABLE XVIII: A comparison of different retain loss settings. w/o GDR refers to using the vanilla forget loss, while w/ KLR indicates the combination with KLR. Overall, $\textsc{PerMU}^{\dagger}$ effectively preserves the model’s utility.
Dataset Forget Rephrased Forget MU↑ FRT↑
Metric RG↓ Pr↓ TR↑ RG↓ Pr↓ TR↑
$\textsc{PerMU}^{\dagger}$ 44.60 47.58 67.25 32.72 37.99 65.96 64.80 1.59
w/o GDR 44.88 47.52 66.44 34.23 39.67 65.50 60.25 1.45
w/ KLR 41.54 45.43 67.45 31.92 38.15 66.44 60.55 1.54
7.9 Different Retain Loss
We investigate the impact of different retain loss functions while keeping other parameters fixed. The experiments are conducted on the TOFU Forget10 dataset using LLaMA2-7B. $\textsc{PerMU}^{\dagger}$ typically employs the forget loss combined with GDR, with a retain weight of $RW=1$. Here, w/o GDR refers to using the vanilla forget loss without any retain loss, while w/ KLR denotes the combination of the vanilla forget loss with KLR, applying the same retain weight. As shown in Tab. XVIII, using the vanilla forget loss achieves strong unlearning performance but may slightly impair model utility. While incorporating KLR can improve model utility, the enhancement is less significant compared to using GDR as the retain loss. Therefore, we primarily adopt GDR as the retain loss in our experiments.
@@ 8 Conclusions
In this paper, we advance LLM-based knowledge forgetting by shifting the focus from superficial forgetting to the more challenging task of implicit knowledge forgetting. To this end, we introduce UGBench, a comprehensive benchmark for evaluating implicit unlearning, and propose PerMU, a novel approach that addresses two core challenges identified by UGBench. Our results highlight the need for more generalised and thorough unlearning in LLMs. While automatic evaluation metrics offer reliable assessments of model performance, we plan to incorporate human evaluations in future work to further enhance the evaluation process. Overall, we believe our contributions offer meaningful insights into effective knowledge unlearning in NLP, with broader implications for related areas such as model editing, knowledge conflict resolution, reasoning, and other diverse fields.

### v3

@@ 5.2 Perturbed Distribution Matching
Building on MSM, we identify the top-KK% most sensitive tokens and inject random noise into their embeddings before feeding them into the model, resulting in the perturbed unlearning sample $x^{\prime}$. We first pass the original unlearning sample $x$ into the unlearning model $f_{\theta_{u}}$ to obtain the clean-run next-token probability distribution $p(y|x)$, where $y=f_{\theta_{u}}(y|x)$ represents the model’s output. Similarly, the corrupted-run next-token probability distribution $p(y|x^{\prime})$ is obtained by feeding the perturbed sample $x^{\prime}$ into $f_{\theta_{u}}$. As illustrated in Fig. 3, the clean-run probability distribution $p(y|x)$ assigns the highest probabilities to tokens that are informative and fact-related. In contrast, for the corrupted-run distribution $p(y|x^{\prime})$, the model fails to recall related facts, resulting in token generation that relies primarily on context or grammar, with top-ranked tokens being fact-unrelated.
@@
1: Identify the top-KK% most sensitive tokens in $x$ using MSM.
2: Introduce random noise to the KK% tokens embeddings to obtain the perturbed sample: $x^{\prime}$.
@@
[…]
@@
6 Evaluation of PerMU
In this section, we evaluate the effectiveness of PerMU and analyze the results. The implementation parameters of PerMU are consistent with those of the baselines. We integrate GDR with $RW=1$, set the percentage of perturbed tokens to $K=0.4$, and maintain perturbation ratio $P=0.4$, coefficient $C=0.1$.
@@ 6.1 Improved Unlearning Capabilities
[…]
@@
[…]
@@ 6.2 Improved Generalisation Capabilities
[…]
@@ 6.3 Discussions
TABLE VIII: Comparison between the average MSM values of subject words and the other words. Subject words exhibit higher MSM values than other tokens, indicating greater sensitivity.
@@ 6.4 Fast Alternative Implementation of PerMU
To reduce the computational overhead of identifying the top-KK% most sensitive tokens for each unlearning sample, we explore a more efficient implementation of PerMU. We begin by computing the MSM to determine the types of tokens to which the model is most sensitive. Experiments are conducted using Phi-1.3B.
@@
We present the experimental results of $\textsc{PerMU}^{\dagger}$. As shown in Tab.IV, $\textsc{PerMU}^{\dagger}$ achieves unlearning performance comparable to existing methods, even surpasses PerMU by up to 17.64% (17.91 $\rightarrow$ 14.75) on Rephrased TOFU Forget01 Probability, and 8.19 (63.09 $\rightarrow$ 68.26) on Rephrased TOFU Forget05 TR. Although $\textsc{PerMU}^{\dagger}$ underperforms PerMU on Harry Potter dataset, it still outperforms baseline methods, achieving relative improvements of up to 15.47% (87.35 $\rightarrow$ 71.88) in Forget ROUGE, 8.69% (65.05 $\rightarrow$ 56.36) in Forget Probability and 15.36% (84.90 $\rightarrow$ 69.54) in Forget F1 when evaluated on LLaMA2-7B, all while maintaining high model utility. These results demonstrate the effectiveness of $\textsc{PerMU}^{\dagger}$ in achieving efficient forgetting with minimal computational overhead.
@@ 7.1 Evaluation on Generation Quality
The impact of machine unlearning on language models is intricate, requiring a thorough and comprehensive evaluation to fully understand its effects. To this end, we perform additional tests to evaluate the generation quality of existing methods. Building on the previous work [38], we introduce the Fluency metric to measure the fluency of the unlearned model’s output sentences. Fluency is measured by the weighted average of bi- and tri-gram entropies [51], defined as $\sum_{k}f(k)\log_{2}f(k)$, where $f(\cdot)$ represents the $n$-gram frequency distribution. A higher Fluency score indicates more informative and diverse text generation. Experiments are conducted on LLaMA2-7B across all subsets.
[…]
@@ 7.4 Computational Overhead
We evaluated the computational overhead of existing methods from two perspectives: training step time and memory usage. The experiments were conducted on the Phi-1.3B model using the TOFU Forget01 dataset, with a batch size of 4 for 5 epochs. We report the total training step time and the maximum memory usage. As shown in Tab. XIV, although PerMU requires more training time, most of the overhead comes from computing the MSM to identify the top-KK% sensitive tokens. Notably, this step can be performed offline prior to training, significantly reducing both time and memory consumption during actual training. Moreover, our proposed fast variant, $\textsc{PerMU}^{\dagger}$, introduces no additional cost in training step time or memory usage, confirming its computational efficiency.
@@ 7.5 Percentage of Perturbed Tokens $K$
In PerMU, we identify the top-$K$ most sensitive tokens and apply perturbations to them. The choice of $K$ directly influences the unlearning effect. We analyze the impact of varying $K$ from 0.1 to 1.0 in increments of 0.1, where each value represents KK% of the tokens to be perturbed, while keeping all other hyper-parameters fixed. As shown in Tab. XIII, increasing $K$ consistently enhances the unlearning ability, with Forget Rouge improved by up to 5.62% (47.19 $\rightarrow$ 41.57) on the TOFU dataset. However, this improvement comes at the cost of reduced model utility. This trade-off is expected, as perturbing more tokens makes it harder for the model to recover and correctly predict the unlearning samples. Although the Forget Retain Trade-off (FRT) score peaks when $K=0.8$ to $0.9$, the performance on retained knowledge drops significantly—up to 5.76% in Rouge (82.04 $\rightarrow$ 76.28) and 3.84% in Probability (88.47 $\rightarrow$ 84.63). Such degradation may negatively impact the model’s utility on preserved knowledge. To strike a balance between forgetting and retention, we set $K=0.4$ in our main experiments, which achieves effective unlearning while preserving most of the model’s capacity to retain relevant knowledge.
@@
TABLE XVI: A comparison of different retain loss settings. w/o GDR refers to using the vanilla forget loss, while w/ KLR indicates the combination with KLR. Overall, $\textsc{PerMU}^{\dagger}$ effectively preserves the model’s utility.
Dataset Forget Rephrased Forget MU↑ FRT↑
Metric RG↓ Pr↓ TR↑ RG↓ Pr↓ TR↑
$\textsc{PerMU}^{\dagger}$ 44.60 47.58 67.25 32.72 37.99 65.96 64.80 1.59
w/o GDR 44.88 47.52 66.44 34.23 39.67 65.50 60.25 1.45
w/ KLR 41.54 45.43 67.45 31.92 38.15 66.44 60.55 1.54
@@ 7.6 Perturbation Ratio $P$
7.7 Different Retain Loss
We investigate the impact of different retain loss functions while keeping other parameters fixed. The experiments are conducted on the TOFU Forget10 dataset using LLaMA2-7B. $\textsc{PerMU}^{\dagger}$ typically employs the forget loss combined with GDR, with a retain weight of $RW=1$. Here, w/o GDR refers to using the vanilla forget loss without any retain loss, while w/ KLR denotes the combination of the vanilla forget loss with KLR, applying the same retain weight. As shown in Tab. XVI, using the vanilla forget loss achieves strong unlearning performance but may slightly impair model utility. While incorporating KLR can improve model utility, the enhancement is less significant compared to using GDR as the retain loss. Therefore, we primarily adopt GDR as the retain loss in our experiments.
TABLE XVII: The impact of the Tuning Coefficient $C$. Experiments are conducted on the TOFU Forget01 dataset using Phi-1.3B. Increasing $C$ results in an improved unlearning effect, but at the cost of decreased model utility. To balance model utility with effective unlearning, we select $C=0.1$ for our experiments.
@@ 7.8 Tuning Coefficient $C$
We investigate the impact of the Tuning Coefficient $C$, which varies from $0.0$ to $1.0$ in increments of $0.1$. The experiments are conducted on the TOFU Forget01 dataset using Phi-1.3B, with all other parameters held constant. As shown in Tab. XVII, the best model utility is achieved when $C=0.0$, where only the corrupted-run probability distribution is used. While this setting maintains high model utility, the unlearning effect is insufficient, as the top-ranked token in the clean-run probability distribution is not fully suppressed. As $C$ increases, the unlearning effect improves, reaching its peak when $C=0.4$. Then it begins to fluctuate as $C$ continues to increase. Correspondingly, model utility decreases with larger values of $C$, which is expected, as higher $C$ values subtract more information from $p(y|x^{\prime})$, potentially disrupting the distribution of irrelevant knowledge. To balance model utility with effective unlearning, we select $C=0.1$ for our experiments.
TABLE XVIII: Experimental results for implementing perturbation at the discrete-token level to subject words, denoted as $\textsc{PerMU}_{dis}$. Surprisingly, $\textsc{PerMU}^{\dagger}_{dis}$ demonstrates a stronger unlearning effect on some datasets, particularly achieving 19.26% in Forget Rouge and 4.64% in Probability on the Forget01 dataset using LLaMA2-7B. These results highlight the adaptability of our approach across different dimensions.
@@
7.9 Discrete-Token Level Perturbation
[…]
@@ 8 Conclusions
In this paper, we advance LLM-based knowledge forgetting by shifting the focus from superficial forgetting to the more challenging task of implicit knowledge forgetting. To this end, we present a comprehensive evaluation and reveal the poor generalization of existing methods in implicit unlearning. To address this limitation, we propose PerMU, which demonstrates strong unlearning performance across multiple real-world scenarios.
In summary, we highlight a promising direction for implicit knowledge forgetting and the need for more thorough and robust unlearning in LLMs.
Limitations
While automatic evaluation metrics provide reliable assessments of model performance, we plan to incorporate human evaluations in future work to further strengthen the evaluation process. Although PerMU does not rely on auxiliary models or classifiers, its perturbation process requires computing token-level gradients, which can incur non-trivial overhead in large-scale settings. To mitigate this, we propose a fast implementation variant with reduced training overhead and cost. The exploration of more efficient methods for handling complex unlearning scenarios that involve reasoning is left to future work.